# Is Grokking a Computational Glass Relaxation?

### v1

@@ 3.3 Wang-Landau Molecular Dynamics (WLMD)
In this section, we will introduce the details of WLMD algorithm and its implementation in NNs. For a physical system made of atoms or particles of other length scales, one can use WLMD to calculate the entropy as a function of the potential energy $S(U)$. This is done by simulating particles with the following equation of motion junghans2014molecular :
@@
Here the Wang-Landau factor $f_{WL}(t)$ quantifies how fast we update the entropy function. In order to ensure the convergence of the WLMD algorithm, usually we make it increase linearly with time to a maximum value $F_{WL}$ and then decay inversely. $\delta$ is a Dirac function, in the simulation we approximated it with a narrow Gaussian function. $U_{0}$ is the potential energy of the system at this moment.
For NNs, we can set the temperature $T_{0}$ and the mass $m_{i}$ to 1, which does not affect the correctness of the WLMD algorithm. The particle coordinates $q_{1},q_{2},...,q_{n}$ are replaced by the NN parameters $c_{1},c_{2},...,c_{n}$, and the entropy we considered is $S(L_{train},A_{test})$. The parameter motion follows
$\frac{d^{2}c_{i}}{dt^{2}}=-\frac{\partial S(L_{train},A_{test},t)}{\partial L_{train}}\frac{\partial L_{train}(c_{1},c_{2},...,c_{n})}{\partial c_{i}}$ (4)
$-\frac{\partial S(L_{train},A_{test},t)}{\partial A_{test}}\frac{\partial A_{test}(c_{1},c_{2},...,c_{n})}{\partial c_{i}}.$
Here, the term $\frac{\partial S}{\partial L_{train}}$ and $\frac{\partial S}{\partial A_{test}}$ can be obtained from numerical differentiation of adjacent bins, and $\frac{\partial L_{train}}{\partial c_{i}}$ is the gradient of $ln(train\ loss)$. $A_{test}$ is a discrete value, to solve its gradient, we use the sigmoid function for approximation. Reflective boundary conditions are used to implement parameter constraints (detailed in Sec. 3.2) for results in Sec. 4.1, and for results in Section 4.2 we restricted the parameter weight norms directly by weight rescaling.
The initialized velocity of the parameters follows the Maxwell–Boltzmann distribution, which is a random Gaussian distribution with mean 00 and variance $k_{B}T/m=1$. In order to maintain constant temperature during the simulation, that is, to conserve the kinetic energy of the parameters, we add the Langevin kinetic term to the velocities:
@@
Here, the definition of $\frac{d^{2}c_{i}}{dt^{2}}$ is given in Equation 4, $c_{f}$ is the friction coefficient, and $v_{rand}$ is the random velocity follows a standard Gaussian distribution. In the NN parameters updating process, by continuously adding entropy to the entropy landscape, $S(L_{train},A_{test},t)$ will eventually converge to the true entropy distribution $S(L_{train},A_{test})$.
When sampling the entropy landscape of grokking, we are interested in the regime with lower training loss, which corresponds to the memorization state and the generalization state. However, the WLMD method tends to explore areas with higher entropy, which are usually areas with higher training loss. In order to prevent the simulation program from wasting time outside the range of our interest, we set an additional entropy bias:
$\mathrm{if}\ L_{train}>L_{0},\ S_{bias}(L_{train})=\beta(L_{train}-L_{0})^{2};\ \mathrm{else}\ S_{bias}(L_{train})=0.$ (6)
If the training loss is outside the range of interest $(L_{train}>L_{0})$ when updating the NN parameters, this entropy bias will provide additional entropy gradients to increase the acceleration to force the NN move towards lower training loss.
@@ 4.1 No entropy barrier in entropy landscape
In thermodynamics, a phase transition is a process in which a material changes from one state of matter to another. Phase transitions are studied with the help of a so-called “order parameter,” which can actually be any quantity that changes significantly during the transition. For example, in studying liquid-to-gas phase transitions, the order parameter can be the mass density, the energy density, or the refractive index. The order parameter is usually discontinuous in first-order phase transitions, but is continuous in second-order and higher-order phase transitions. When training an NN with grokking, it quickly goes from underfitting to memorization, and then takes a long time to reach generalization. If the NN is regarded as a physical system, then the transition from the memorization state to the generalization state appears to be a phase transition. If so, the test accuracy is a natural choice of the order parameter.
@@
A previous study suggests that grokking is a first-order phase transition process from memorization to generalization rubin2023grokking . They used the Langevin dynamic optimizer to train two two-layer nonlinear teacher-student models and found that the memorization and generalization states of the NN were located at two free-energy minima, thus believing that grokking is a first-order phase transition that requires crossing a free-energy barrier. We directly examined this statement by computing the entropy landscape as a function of training loss and test accuracy to search for such a barrier. We examine three algorithmic tasks on a one-layer transformer model. The three tasks range from easy to learn, to difficult to learn, and to impossible to learn. The training results are also given in the Appendix Figure 5, where grokking is observed for the first two tasks. Using the WLMD method, we sample the entropy landscape of training loss and test accuracy as presented in Figure 1. For the first two tasks that exhibit grokking, the training trajectory is roughly “L” shaped. In the vertical part of the trajectory, i.e., from the memorization state (lower left part) to the generalization state (upper left part), we observed no entropy barrier. Since there is neither an entropy barrier nor an energy (train loss) barrier, our results disproves the view that grokking requires crossing a free-energy barrier. In addition, we find that the equilibrium (maximum entropy for a given training loss) curve in the low train loss region is continuous, which also denies the first-order phase transition, as such transitions require discontinuity of the order parameter (test accuracy). For the third unlearnable task, the equilibrium state also fails to generalize, although some states present a higher test accuracy (about 20%) than random selection (about 1.5%).
[…]
@@
We also found a huge gap between the AdamW training curve and the equilibrium state curve. We define this significant advantage of the equilibrium state over the training curve as a high-entropy advantage under grokking. This result shows that grokking is actually a kinetic process rather than a phase change in the thermodynamic sense: The NN is far away from the equilibrium state during the process of minimizing the training loss, and finally reaches the equilibrium state, which generalizes well, after a long time. This process is similar to glass formation and relaxation: When the liquid is cooled rapidly, the molecules do not have enough time to rearrange themselves into a crystal or crystal-like structure, forming a disordered glassy state (non-equilibrium state), which then gradually evolves toward a better equilibrated state through slow structural relaxation.
@@ 4.2 Neural network without grokking does not exhibit glassy behavior
The analogy between grokking and glass relaxation can be further verified by examining the entropy landscape after eliminating grokking. According to our conjecture, just as a slowly cooled liquid will maintain thermodynamic equilibrium and will not enter the non-equilibrium glassy state, an NN without grokking will be trained along the equilibrium curve as the training loss decreases, steadily reaching the generalization state. The study by Liu et al. liu2022omnigrok proposed that grokking can be eliminated by limiting the weight norm of the NN parameters in an appropriate range. In this study, we kept the weight norm of the parameters at 30, and the training results presented in Appendix Figure 6 show that grokking can be greatly eliminated under this condition. Here we consider three different Transformer models with different widths (model dimensions).
@@
Figure 2 shows that after controlling the weight norm at 30, the AdamW training curve in the entropy landscape is very close but still slightly below the equilibrium state. Moreover, under this restriction the entropy landscape cannot explore serious memorization states, which is the computational glassy state in our analogy. This result further illustrates the correspondence between glass and grokking, which is consistent with our conjecture. In addition, we note that the high-entropy advantage still exists even in the network where grokking is eliminated. This agrees with the work of Yang et al. yang2025high that high-entropy advantages are prevalent in various deep learning tasks, and the connection between equilibrium and generalization is strong and extensive.
@@ 5 WanD: Eliminating grokking without regularization
$\frac{d^{2}\theta(t)}{dt^{2}}=-\frac{\partial S(L_{train}(t))}{\partial L_{train}(\theta(t))}\cdot\frac{\partial L_{train}(\theta(t))}{\partial\theta(t)}+\eta(t)-\lambda\frac{d\theta(t)}{dt},$ (7)
where $\theta(t)$ is a tensor of the NN parameters in time $t$, and $L_{train}$ is the logarithm of the cross-entropy training loss. $\eta(t)$ is a noise given by $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\lambda(2-\lambda)k_{B}T\delta_{ij}\delta(t-t^{\prime})$ , $k_{B}T=1$ is a temperature term, and $\lambda$ is a friction coefficient. When sampling the entropy landscape using WLMD, we obtain the corresponding training loss and test accuracy of the NN to locate where to add entropy. In designing the training optimizer, we cannot use the test data, so the training process can be regarded as sampling the entropy $S(L_{train})$ in the dimension of the training loss alone. In addition, we set up an appropriate entropy bias to accelerate the NN’s exploration at low training loss area.
The training results for modular addition task using WanD or AdamW are shown in Figure 3. Both methods have comparable performance, achieving generalization after 2,000 optimization steps. During the training process of WanD, the weight norm of the NN parameters continued to increase, eventually reaching about 175, indicating that the NN can generalize at high weight norm. This agrees with the work of Kumar et al. kumar2023grokking that regularization is not necessary for grokking, but in a more strictly-defined context as we do not scale NN’s output, which could be taken as effectively changing the weight norm.
Figure 3: Using (a)(b)(c) WanD and (d) AdamW to train a one layer Transformer on modular addition ($p=67$) task. (a) Train and test accuracy over time. The grokking is largely eliminated while training with WanD. (b) Weight norm over time. (c) Test accuracy over weight norm. These together demonstrate that the model generalized at high weight norm. (d) Train and test accuracy over time. Both WanD and AdamW using the same learning rate.
@@ 6 Conclusions and Discussion
In this work, we studied grokking in the modular arithmetic tasks on a one-layer transformer, and found entropy as an excellent indicator of generalization. Our results agree with previous findings feng2023activity ; neyshabur2017exploring that flatness should be combined with the weight norm in order to reflect generalizability. This agreement can be understood by considering a set of adjacent points representing parameter configurations with the same generalizability in the parameter space. Since weight scaling can be performed without changing the generalization performance dinh2017sharp , this set forms a cone starting from the origin of the parameter space and extending to infinity. Within this cone, the flatness of the solution depends on the weight norm. Our entropy measure, defined as the logarithm of the volume of parameter space occupied by this set of solutions within some cutoff, is roughly a measure of the ’conical degree’, which is a quantity combining both flatness and weight norm. Our findings show that the generalization set has a larger conical degree than the memorization set, supporting these previous findings feng2023activity ; neyshabur2017exploring . Similar correlation between entropy and generalizability has been found for other neural networks and datasets yang2025high .
By sampling the entropy landscape of the network on training loss and test accuracy, we found that i). there is no entropy barrier between the memorization state and the generalization state, which means that the memorization state is not in an equilibrium or meta-stable state. Why does grokking take so long to happen even without crossing free energy barriers? In the last part of this section, we postulate that the reason is the formation of glass states. ii). The equilibrium test accuracy, as a function of the training loss, appears continuous. This strongly suggests that that grokking is not a first-order phase transition as argued by previous work rubin2023grokking , because a first-order phase transition implies discontinuity of physical quantities. Our results also indicate that the equilibrium state in the entropy landscape has a significant advantage over the training curve, which we refer to it as the high-entropy advantage under grokking. Such a high-entropy advantage is weakened but still apparent when grokking is eliminated via fixing weight norm. This demonstrates a correlation between equilibrium state and NNs’ generalization, consistent with the conclusion of Yang et al.yang2025high . In addition, we designed a toy optimizer based on the Wang-Landau molecular dynamics method, WanD, which encourages exploration of the high entropy state for NNs’ training. WanD successfully achieves generalization with a similar time efficiency as AdamW but avoids grokking. The generalization states found by WanD have a much higher weight norm compared with the one needed to eliminate grokking in traditional training liu2022omnigrok . This further supports the conclusions of Kumar et al. kumar2023grokking that grokking is not solely due to weight decay as some theory proposed before liu2022omnigrok in a strictly-defined context.

### v2

@@ 3.3 Wang-Landau Molecular Dynamics (WLMD)
In this section, we will introduce the details of WLMD algorithm and its implementation in NNs. The key advantage WLMD is its ability to avoid becoming trapped junghans2014molecular . In contrast to standard samplers, which tend to oversample the high-entropy regions, WLMD achieves uniform sampling across all regions of the entropy diagram upon convergence kim2006statistical . This is accomplished by maintaining a running estimate of the entropy landscape, which introduces an ‘entropy gradient’ that drives the system away from previously visited states. The convergence of the WLMD is established in fort2015convergence . For a physical system made of atoms or particles of other length scales, one can use WLMD to calculate the entropy as a function of the potential energy $S(U)$. This is done by simulating particles with the following equation of motion junghans2014molecular :
@@
Here the Wang-Landau factor $f_{WL}(t)$ quantifies how fast we update the entropy function. In order to ensure the convergence of the WLMD algorithm, usually we make it increase linearly with time to a maximum value $F_{WL}$ and then decay inversely. $\delta$ is a Dirac function, and we approximated it with a narrow Gaussian function in the simulation. $U_{0}$ is the potential energy of the system at this moment.
For NNs, we can set the temperature $T_{0}$ and the mass $m_{i}$ to 1, which does not affect the correctness of the WLMD algorithm. The particle coordinates $q_{1},q_{2},...,q_{n}$ are replaced by the NN parameters $c_{1},c_{2},...,c_{n}$, and the entropy we considered is $S(x_{WL},y_{WL})$. We choose $ln(L_{train})$ as the energy in a broad sense $x_{WL}$, and $A_{test}$ as the order parameter $y_{WL}$. The parameter motion follows
$\frac{d^{2}c_{i}}{dt^{2}}=-\frac{\partial S(x_{WL},y_{WL},t)}{\partial c_{i}}=-\frac{\partial S(ln(L_{train}),A_{test},t)}{\partial ln(L_{train})}\frac{\partial ln(L_{train}(c_{1},c_{2},...,c_{n}))}{\partial c_{i}}$ (4)
$-\frac{\partial Sln(L_{train}),A_{test},t)}{\partial A_{test}}\frac{\partial A_{test}(c_{1},c_{2},...,c_{n})}{\partial c_{i}}.$
The term $\frac{\partial S}{\partial ln(L_{train})}$ and $\frac{\partial S}{\partial A_{test}}$ can be computed via numerical differentiation between adjacent bins, while $\frac{\partial ln(L_{train})}{\partial c_{i}}$ is the gradient of ln(trainloss)ln(train\ loss) over parameters. Since $A_{test}$ is discrete, we approximate its gradient using the sigmoid function, following the approach in yang2025high . Further details are also provided in Appendix A.3. Reflective boundary conditions are used to implement parameter constraints for results in Sec. 4.1, and for results in Section 4.2 we restricted the parameter weight norms directly by weight rescaling (both detailed in Sec. 3.2).
The initialized velocity of the parameters follows the Maxwell–Boltzmann distribution, which is a random Gaussian distribution with mean 0 and variance $k_{B}T/m=1$. In order to maintain constant temperature during the simulation, that is, to conserve the kinetic energy of the parameters, we add the Langevin kinetic term to the velocities:
@@
Here, the definition of $\frac{d^{2}c_{i}}{dt^{2}}$ is given in Equation 4, $c_{f}$ is the friction coefficient, and $v_{rand}$ is the random velocity follows a standard Gaussian distribution. In the NN parameters updating process, by continuously adding entropy to the entropy landscape, $S(x_{WL},y_{WL},t)$ will eventually converge to the true entropy distribution $S(x_{WL},y_{WL})$.
Due to space limitations, more details of WLMD and the complete algorithm are described in the Appendix A.3. To the best of our knowledge, WLMD is currently the most efficient entropy sampling algorithm, as existing alternatives are limited to models with only hundreds of parameters yang2025high . The computational complexity of WLMD scales linearly with the number of NN parameters, and the entropy landscape sampling can be parallelized across multiple GPUs. This enables entropy sampling for large-scale models, such as transformers with hundreds of thousands of parameters.
@@ 4.1 No entropy barrier in entropy landscape
In thermodynamics, a phase transition is marked by a significant change in an order parameter—a macroscopic, measurable quantity like density during a liquid–gas transition. By analogy, the learning dynamics of NNs exhibiting grokking appears to mimic the characteristics of a phase transition: starting with a rapid progress from underfitting to memorization (low test accuracy), the NN undergoes an abrupt qualitative shift to generalization (high test accuracy) after a prolonged training period. Given this, we identify the test accuracy as the order parameter for characterizing grokking. Just as a physical order parameter changes distinctly between material states, the definitive increase in test accuracy can track the NN’s transition into the generalization regime, making it a natural and compelling macroscopic indicator of this critical learning phenomenon.
@@
A previous study suggests that grokking is a first-order phase transition process from memorization to generalization rubin2023grokking . They used the Langevin dynamic optimizer to train two two-layer nonlinear teacher-student models and found that the memorization and generalization states of the NN were located at two free-energy minima, thus believing that grokking is a first-order phase transition that requires crossing a free-energy barrier. We directly examined this statement by computing the entropy landscape as a function of training loss and test accuracy to search for such a barrier. We examine three algorithmic tasks on a one-layer transformer model. The three tasks range from easy to learn, to difficult to learn, and to impossible to learn. The training results are also given in the Appendix Figure 5, where grokking is observed for the first two tasks. Using the WLMD method, we sample the entropy landscape of training loss and test accuracy as presented in Figure 1. The three entropy landscapes are the results after 60M, 80M, and 20M WLMD update epochs respectively, where these durations were chosen as the points where the entropy map exhibited no substantial further change. We ran 8 WLMD processes on 4 GeForce RTX 4090 GPUs, requiring approximately 100 GPU hours per 1M epochs. Additional computational details are provided in Appendix A.2.
For the first two tasks that exhibit grokking, the training trajectory is roughly “L” shaped. In the vertical part of the trajectory, i.e., from the memorization state (lower left part) to the generalization state (upper left part), we observed no entropy barrier. Since there is neither an entropy barrier nor an energy (train loss) barrier, our results disproves the view that grokking requires crossing a free-energy barrier. In addition, we find that the equilibrium (maximum entropy for a given training loss) curve in the low train loss region is continuous, which also denies the first-order phase transition, as such transitions require discontinuity of the order parameter (test accuracy). For the third unlearnable task, the equilibrium state also fails to generalize, although some states present a higher test accuracy (about 20%) than random selection (about 1.5%).
Figure 1: Entropy landscape of three different arithmetic tasks : (a) x+ymod 67x+y\ mod\ 67 is easy to learn, max entropy state can generalize better than training line. (b) $x^{2}+y\ mod\ 67$ is hard to learn, the generalization advantage of the max entropy state over the training results under the same training loss is very significant. (c) $x^{3}+xy^{2}+y\ mod\ 67$ cannot be learned by training, max entropy state also can not generalize.
@@
We also found a huge gap between the AdamW training curve and the equilibrium state curve. We define this significant advantage of the equilibrium state over the training curve as the high-entropy advantage under grokking. This result shows that grokking is actually a kinetic process rather than a phase transition in the thermodynamic sense: The NN is far from the equilibrium state during the initial part of the training process, and finally reaches the equilibrium state, which generalizes well, after a long time. This process is similar to glass formation and relaxation: When the liquid is cooled rapidly, the molecules do not have enough time to rearrange themselves into a crystal or crystal-like structure, forming a disordered glassy state (non-equilibrium state), which then gradually evolves toward a better equilibrated state through slow structural relaxation.
@@ 4.2 Neural network without grokking does not exhibit glassy behavior
The analogy between grokking and glass relaxation can be further verified by examining the entropy landscape after eliminating grokking. According to our conjecture, just as a slowly cooled liquid will maintain thermodynamic equilibrium and will not enter the non-equilibrium glassy state, an NN without grokking will be trained along the equilibrium curve as the training loss decreases, steadily reaching the generalization state. The study by Liu et al. liu2022omnigrok suggested that grokking can be eliminated by limiting the weight norm of the NN parameters in an appropriate range. Following this insight, we implemented a strict weight norm constraint in our study: after each iteration step, we compute the weight norm of the neural network parameters and scale them proportionally to maintain a fixed norm of 30. The training results, presented in Appendix Figure 6, show that grokking can be greatly eliminated under this condition. We examine three transformer models with varying widths (model dimensions). For each entropy sampling task, we use 4 NVIDIA H100 GPUs to run 8 processes, totaling 60M epochs and consume nearly 4,000 GPU hours. Further computational details are available in Appendix A.2.
@@
Figure 2 shows that after controlling the weight norm at 30, the AdamW training curve in the entropy landscape is very close but still slightly below the equilibrium state. Moreover, under this restriction the entropy landscape cannot explore severe memorization states, which is the computational glassy state in our analogy. This result further illustrates the correspondence between glass and grokking. In addition, we note that the high-entropy advantage still exists even in the network where grokking is eliminated. This agrees with the work of Yang et al. yang2025high that high-entropy advantages are prevalent in various deep learning tasks and there is a robust connection between equilibrium and generalization.
@@ 5 WanD: Eliminating grokking without regularization
$\frac{d^{2}\theta(t)}{dt^{2}}=-\frac{\partial S(ln(L_{train}(\theta(t))))}{\partial ln(L_{train}(\theta(t)))}\cdot\frac{\partial ln(L_{train}(\theta(t)))}{\partial\theta(t)}+\eta(t)-\lambda\frac{d\theta(t)}{dt},$ (6)
where $\theta(t)$ is a tensor of the NN parameters in time $t$, and $ln(L_{train})$ is the logarithm of the cross-entropy training loss. $\eta(t)$ is a noise given by $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\lambda(2-\lambda)k_{B}T\delta_{ij}\delta(t-t^{\prime})$ , $k_{B}T=1$ is a temperature term, and $\lambda$ is a friction coefficient. When sampling the entropy landscape using WLMD, we obtain the corresponding training loss and test accuracy of the NN to locate where to add entropy. For the WanD optimizer, its training process samples the entropy landscape $S(ln(L_{train}))$ based solely on training loss, due to the fundamental constraint that it can only access training data. In addition, we set up an appropriate entropy bias to accelerate the NN’s exploration at low training loss area. The overall WLMD entropy sampling procedure is summarized in Algorithm 1.
Algorithm 1 WanD Optimizer
[…]
Transformer model $T$ reach generalization: $Accuracy(T(X_{test}),y_{test})\approx 100\%$
The initial entropy: $S\leftarrow S_{bias}$
for each optimize epoch do
  Compute the logarithm of the train loss:
  $ln(L_{train})=ln(CrossEntropyLoss(T(X_{train}),y_{train}))$;
  Compute entropy gradient $x_{grad}=\frac{dS(x_{WL})}{dx_{WL}}$
  Compute $\frac{d^{2}c_{i}}{dt^{2}}=-\frac{\partial ln(L_{train})}{\partial c_{i}}x_{grad}$
  for each parameter $c_{i}$ do
   Update the parameter velocity: $\frac{dc_{i}}{dt}\leftarrow\frac{dc_{i}}{dt}(1-c_{f})+\frac{d^{2}c_{i}}{dt^{2}}\Delta t+v_{Langevin}$
   Update the parameter: $c_{i}\leftarrow c_{i}+\frac{dc_{i}}{dt}\Delta t$
  end for
  Update the Wang-Landau factor by Equation 3: $f_{WL}\leftarrow F_{WL}min\{t/t_{WL},t_{WL}/t\}$
  Update the entropy landscape by Equation 3: $S(x_{WL})\leftarrow S(x_{WL})+f_{WL}\delta(x-x_{WL})$
[…]
end for
The training results for modular addition task using WanD or AdamW are shown in Figure 3. Both methods have comparable performance, achieving generalization after $\sim 2,000$ optimization steps. WanD has a shorter running time, taking about 50 seconds per 1K steps compared to roughly 72 seconds for AdamW on a single GeForce RTX 4090 GPU. During the training process of WanD, the weight norm of the NN parameters continued to increase, eventually reaching about 175, indicating that the NN can generalize at high weight norm. This agrees with the work of Kumar et al. kumar2023grokking that regularization is not necessary for grokking, but in a more strictly-defined context as we do not scale NN’s output, which could be taken as effectively changing the weight norm. We present additional training results for WanD and AdamW in the Appendix B.3 on the same task but eliminating grokking by increasing the training dataset split fraction.
Figure 3: Using (a)(b)(c) WanD and (d) AdamW to train a one layer Transformer on modular addition ($p=67$) task. (a) Train and test accuracy over time. The grokking is largely eliminated while training with WanD. (b) Weight norm over time. (c) Test accuracy over weight norm. These (a, b, c) together demonstrate that the model can generalize at high weight norm. (d) Train and test accuracy over time. Both WanD and AdamW use the same learning rate.
@@ 6 Conclusions and Discussion
6.1 Conclusions
In this work, we investigated grokking in the modular arithmetic tasks using a one-layer transformer, and found that Boltzmann entropy (logarithm of solution volume) as an excellent indicator of generalization. By sampling the entropy landscape of the network on training loss and test accuracy, we found that i). there is no entropy barrier between the memorization state and the generalization state, which means that the memorization state is not in an equilibrium or meta-stable state. Why does grokking take so long to happen even without crossing free energy barriers? In the last part of this section, we postulate that the reason is the formation of glass states. ii). The equilibrium test accuracy, as a function of the training loss, appears continuous. This strongly suggests that that grokking is not a first-order phase transition as argued by previous work rubin2023grokking , because a first-order phase transition implies discontinuity of physical quantities. Our results also indicate that the equilibrium state in the entropy landscape has a significant advantage over the training curve, which we refer to it as the high-entropy advantage under grokking. Such a high-entropy advantage is weakened but still apparent when grokking is eliminated via fixing weight norm. This demonstrates a correlation between equilibrium state and NNs’ generalization, consistent with the conclusion of Yang et al.yang2025high . In addition, we designed a toy optimizer based on the Wang-Landau molecular dynamics method, WanD, which encourages exploration of the high entropy state for NNs’ training. WanD successfully achieves generalization with a similar time efficiency as AdamW but avoids grokking. The generalization states found by WanD have a much higher weight norm compared with the one needed to eliminate grokking in traditional training liu2022omnigrok . This further supports the conclusions of Kumar et al. kumar2023grokking that grokking is not solely due to weight decay as some theory proposed before liu2022omnigrok in a strictly-defined context.
6.2 Discussion